• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 00:04:02 on October 27, 1998

    Entry number 3791

    keyword=Beam Time Accounting
    Hall A

    10/26/1998 1600-2400
    Experiment #: E94-010
    Shift Leader: K. Slifer
    Shift Workers: M.Finn, T.Averett, P.Djawotho

    16001700180019002000210022002300Total Time
    Acceptable Beam
    in Use
    1 1 25 1 55 12 55 1 6.45
    Beam Available
    but Not in Use
    0 0 35 0 0 45 5 0 1.41666666666667
    Beam Not Available
    or Unacceptable
    0 0 0 0 5 3 0 0 0.133333333333333
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    1 1 25 1 55 12 55 1 6.45
    Planned Config-
    uration Change
    0 0 35 0 0 45 5 0 1.41666666666667
    Unplanned
    Experiment Down
    0 0 0 0 5 3 0 00.133333333333333
    Hall not
    ready reasons
    Target 0 0 35 0 0 0 0 0 0.583333333333333
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: 16:00-18:00 (2.796Gev perp) 18:00-18:35 Target NMR, Pedestal run 18:35-20:55 2.796 Gev perp 20:55-21:03 rf trip -->no beam to hall 21:03-21:15 beam restored 2.796 Gev perp 21:15 completed 2.796 |_ (30.8Mevents) began momentum change, target NMR 22:05-22:39 magnets settled. Raster tests, target beam warmup 22:39 began production data on \\2.5804Gev

    Data Entry by: slifer